# Materials Map Integrating Experimental and Computational Data through Graph-Based Machine Learning for Enhanced Materials Discovery

### v2

@@ IV.3 Finding proper materials with material maps
Finding optimal materials is extremely important not only in basic research but also in industrial applications, but is always challenging due to the enormous variety of candidate materials. High-throughput screening of materials using computational databases and machine learning is a powerful tool that can suggest promising materials. However, experimental materials researchers frequently face difficulties in developing suggested new materials, due to limited access to raw materials, equipment, facilities, and experimental expertise. Therefore, it is more important to explore promising materials among those that experimental researchers can realistically fabricate.
@@
The material maps developed in this study have great potential to make materials development more efficient because materials with similar structures are likely to be synthesized and characterized using similar methods. For example, researchers working on high-entropy materials, which have significantly complex compositions and structures, require different knowledge and experimental equipment compared to those investigating ordinary materials. The interactive graph, shown in Figure 8, is a powerful tool for targeting specific material properties by referencing aspects such as composition, data names, and other material properties, enabling multifaceted and detailed data exploration. Using Python libraries such as Plotly, interactive graphs can be easily created, and when saved as HTML files, the data can be accessed at any time. The corresponding HTML file is included in the SI. In conclusion, the material maps, which organize materials data based on structural complexity, can assist researchers in screening candidate materials for experimental studies.
IV.4 Prospects for future studies
@@
In our study, we integrate computational and experimental data by predicting experimental values from compositions [25] by machine learning with experimental dataset of StarryData2 [24]. However, another approach—predicting stable structure from composition—is also conceivable. One should select the method based on the accuracies of the machine learning models, opting for the approach that achieves higher predictive accuracy. In this study, we employed our approach due to the high prediction accuracy of our model ( 0.9), predicting experimental values from compositions, and the technical difficulty in the prediction of the stable structure by composition [27]. Moreover, our study focused on the materials containing six selected elements (Sb, Te, Sn, Se, Bi, S), which demonstrated predictive capability for thermoelectric performance through the values [25]. Extending this model to encompass a wider range of elements and other physical parameters is left for future studies.
V Conclusion
@@
VI Acknowledgments
@@
VII Supplemental information
VII.1 Data Preparation and Machine Learning
VII.1.1 Cleaning of Data in StarryData2
@@
VII.1.2 Machine Learning Modeling of Experimental Data
@@
VII.1.3 Machine Learning Modeling of Calculation Data
@@
VII.2 Correlation analysis of parameters
@@
VII.3 Dimensionality Reduction for Materials Map Construction
In this study, we generate material maps by dimensional reduction using the t-SNE algorithm [9], which is implemented using the scikit-learn framework, as in MDL. The default hyperparameters in MDL, which are a perplexity of 50 and a learning rate of 300, were used for the dimensional reduction by t-SNE.
VII.4 Hardware and Software Environment

### v3

@@ 4.3 Finding proper materials with material maps
Finding optimal materials is extremely important not only in basic research but also in industrial applications, but is always challenging due to the enormous variety of candidate materials. High-throughput screening of materials using computational databases and machine learning is a powerful tool that can suggest promising materials. However, experimental researchers of materials frequently face difficulties in developing suggested new materials, due to limited access to raw materials, equipment, facilities, and experimental expertise. Therefore, it is more important to explore promising materials among those that experimental researchers can realistically fabricate.
@@
The material maps developed in this study have great potential to make materials development more efficient because materials with similar structures are likely to be synthesized and characterized using similar methods. For example, researchers working on high-entropy materials, which have significantly complex compositions and structures, require different knowledge and experimental equipment compared to those investigating ordinary materials. The interactive graph, shown in Figure 8, is a powerful tool for targeting specific material properties by referencing aspects such as composition, data names, and other material properties, enabling multifaceted and detailed data exploration. Using Python libraries such as Plotly, interactive graphs can be easily created, and when saved as HTML files, the data can be accessed at any time. The corresponding HTML file is included in summplementary data. In conclusion, the material maps, which organize materials data based on structural complexity, can assist researchers in screening candidate materials for experimental studies.
4.4 Prospects for future studies
@@
In our study, we integrate computational and experimental data by predicting experimental values from compositions [32] by machine learning with experimental dataset of StarryData2 [31]. However, another approach—predicting stable structure from composition—is also conceivable. One should select the method based on the accuracies of the machine learning models, opting for the approach that achieves higher predictive accuracy. In this study, we employed our approach due to the high prediction accuracy of our model ( 0.9), predicting experimental values from compositions, and the technical difficulty in the prediction of the stable structure by composition [34].
Our study focused on the materials containing six selected elements (Sb, Te, Sn, Se, Bi, S), which demonstrated predictive capability for thermoelectric performance through the values [32]. Extending this model to encompass a wider range of elements and other physical parameters is left for future studies.
5 Conclusion
@@
6 Author contributions
Li H and Takaaki T supervised the project; Yusuke H conceived the idea; Jia X performed the data preprocessing; Yusuke H performed data analysis and wrote the paper; all authors discussed the results; all listed authors agree to all manuscript contents, the author list and its order, and the author contribution statements.
7 Conflict of interest
The authors declare that they have no conflict of interest.
8 Data availability
The corresponding raw data files as well as further data that support the findings of this work are available from the corresponding authors upon reasonable request.
9 Code availability
The code used for data analysis and simulated results is available from the corresponding authors upon reasonable request.
10 Acknowledgments
@@
Appendix A Data Preparation and Machine Learning
A.1 Cleaning of Data in StarryData2
@@
A.2 Machine Learning Modeling of Experimental Data
@@
A.3 Machine Learning Modeling of Calculation Data
@@
Appendix B Correlation analysis of parameters
@@
Appendix C Dimensionality Reduction for Materials Map Construction
In this study, we generate material maps by dimensional reduction using the t-SNE algorithm, which is implemented using the scikit-learn framework, as in MDL. The default hyperparameters in MDL, which are a perplexity of 50 and a learning rate of 300, were used for the dimensional reduction by t-SNE.
Appendix D Hardware and Software Environment